# Reinforced Prompt Personalization for Recommendation with Large Language Models

### v1

@@ 3.4. Actor-Critic Based Architecture of the Multi-agent and Reward Function
To personalize each pattern individually with the corresponding agent, the Multi-Agent Reinforcement Learning (MARL) is structured with four Actor-Critics (Konda and Tsitsiklis, 1999) under CTDE (Lowe et al., 2017; Sunehag et al., 2018), where each agent has independent parameters but has access to the global state, as shown in Figure 2. The Actor and Critic of each agent are two-layer fully connected neural networks parameterized by $\theta_{g}^{(k)}$ and $\theta_{h}^{(k)}$, denoted as $g^{(k)}(\cdot)$ and $h^{(k)}(\cdot)$, respectively. Specifically, each Actor $g^{(k)}(\cdot)$ makes policy to obtain action $a_{t}^{(k)}$ and its corresponding probability $prob_{t}^{(k)}$ based on the state $\mathbf{s}_{t}$, while each Critic $h^{(k)}(\cdot)$ evaluates the value $v_{t}^{(k)}$ of state $\mathbf{s}_{t}$. Formally, we have:
@@
Since the aim is to boost the recommendation ability of LLMs by personalizing instance-wise prompts, we leverage Normalized Discounted Cumulative Gain (NDCG) directly as the reward, which is the performance metric for ranking tasks. Here we adopt $NDCG@M$ as the reward evaluated according to the answer of LLMs, where $M$ is the number of candidate items. Formally, given the recommendation output $o_{t}$ at step $t$, we can define the reward $r_{t}$ as below:
@@
(8) $L_{c}^{(k)}$ $=\frac{1}{N}\sum_{t}(\hat{R}_{t}-v_{t}^{(k)}),$
(9) $L_{a}^{(k)}$ $=\frac{1}{N}\sum_{t}\mathrm{log}(prob_{t}^{(k)}(\hat{R}_{t}-v_{t}^{(k)})),$
@@ 3.5. Algorithm
The algorithm for our framework is as follows.
@@
Algorithm 1 Reinforced Prompt Personalization
@@ 4. Experiments
We evaluate the effectiveness of RPP/RPP+ on LLaMa2-7B-chat (Touvron et al., 2023), comparing it against various baselines including traditional recommender models, few-shot methods, and prompt-based methods. Besides, we extend the evaluation on diverse frozen pre-trained LLMs to measure the RPP/RPP+’s generalization capability. To gain insights into the impact of different patterns in prompts, we conduct an ablation study on personalizing each pattern. Furthermore, we investigate the sensitivity of RPP/RPP+ to hyper-parameters. The case study is employed to provide a tangible illustration of the significance of personalizing prompts in enhancing LLMs’ recommendation ability.
@@ 4.1.1. Datasets
Table 2. Performance comparison of our framework on LLaMa2-7B-chat with baselines across diverse datasets. The results of the best-performing baseline are denoted with an underline. The results of our method which outperform other prompt-based methods are indicated with $\uparrow$, while those that surpass the best-performing baseline are highlighted in red. The performance improvement of RPP/RPP+ compared with the best-performing prompt-based method is listed in the “improvement”.
@@
Improvement 0.78% 0.49 0.45 0.79 0.51 0.49 0.88 0.54 0.48
4.1.2. Baselines
We employ several baseline models for comparison, categorizing them into three types: (1) traditional recommender models, including Pop, BPRMF (Rendle et al., 2009) and SASRec (Kang and McAuley, 2018); (2) few-shot models, such as BM25 (Robertson and Zaragoza, 2009), VQ-Rec (Hou et al., 2023) and UniSRec (Hou et al., 2022); (3) prompt optimization methods, including manual prompts (Petroni et al., 2019; Schick and Schütze, 2021) and enumeration that is a kind of heuristic prompts (Jiang et al., 2020; Gao et al., 2021; Prasad et al., 2023).
Traditional recommender models: Pop relies on item popularity for recommendations but overlooks personalization. BPRMF (Rendle et al., 2009) combines Bayesian methods and matrix factorization techniques to decompose the user-item interaction matrix into latent factor matrices. SASRec (Kang and McAuley, 2018) leverages a self-attention encoder to comprehensively analyze the long-term dependencies present in user behaviors. And it predicts subsequent interactions through a feed-forward network. They are evaluated on the candidate sets after training on the training dataset with all interactions following the setting of LLMRank (Hou et al., 2024)
Prompt-based methods: Manual prompts (Petroni et al., 2019; Schick and Schütze, 2021) are prompts designed by humans to guide LLMs’ behavior, relying on human expertise and domain knowledge. Here we meticulously craft a task-wise prompt for the ranking tasks, encompassing the key patterns within a prompt. Enumeration (Prasad et al., 2023) entails automatically searching within several possible prompts and selecting the best one. Here we design diverse prompts with variations for the ranking tasks, adopting enumeration followed by selection.
4.1.3. Frozen Pre-trained LLMs
To evaluate the generalization ability of RPP/RPP+, we extend the experiments on diverse frozen pre-trained LLMs, including LLaMa2(Touvron et al., 2023), ChatGPT, and Alpaca (Taori et al., 2023). They serve as the representative LLMs of three types respectively: the open-source LLM, which can be inferred using pre-trained parameters; the black-box LLM, which solely offers an accessible API; and the fine-tuned LLM, tailored specifically for the recommendation task. LLaMa2 (Touvron et al., 2023) is an open-source LLM released by Meta whose performance has been widely recognized. Here we select the version of LLaMa2-7B-chat as the frozen LLM. ChatGPT is a black-box LLM, which is excellent across a wide array of tasks. We conduct experiments by utilizing the API of GPT-3.5-turbo from OpenAI and set the temperature to 0.2. Alpaca (Taori et al., 2023) is a lightweight instruction-tuning strategy based on LLaMa. We select LLama2- 7B as the backbone and fine-tune Alpaca with Lora (Hu et al., 2022), which adds trainable rank decomposition matrices to each layer of the transformer architecture in the frozen LLM (Bao et al., 2023), enabling easier and more efficient fine-tuning. We transform the recommendation data for ranking tasks into instruction-tuning data, ensuring that LLMs grasp the specific nuances of the ranking task. The format of the data for instruction tuning can be articulated as:
[…]
4.2. Implementation Details
We configure the initial length of interaction history $l_{0}$ as 1 and the number of candidate items to be $M=10$. In our experiments, we set 3, 9, and 5 sentence choices for “role-playing”, “reasoning guidance”, and “output format” patterns in the action spaces respectively. Candidate items are randomly selected from the training dataset and include a ground-truth item sourced from the test dataset. Therefore, for ranking tasks, the HitRatio@10 metric is always 1, rendering this metric meaningless. We set the $\gamma$ in the reward as 0.95. To reduce the randomness of LLMs, we set the temperature as 0.2, since a lower temperature favors more conservative and deterministic outputs. Besides, we report the average performance with corresponding standard deviations of prompt-based methods and RPP/RPP+, based on at least five repeated runs. Traditional recommender models and few-shot-based methods are executed on an NVIDIA RTX 3090Ti, while prompt-based methods and our framework, based on LLMs, run on an NVIDIA A100. During the training phase, we randomly select 200 users as training examples to optimize the multi-agent strategy employed in RPP/RPP+. We terminate the iterative training process if the highest NDCG@10 does not update over 7 iterations or if the total iterations exceed 15. Subsequently, during the testing phase, the multi-agent, now equipped with an optimized policy, undergoes evaluation on a separate set of 100/200 test examples.
@@ 4.2. Implementation Details
Table 3. Performance of RPP and RPP+ on diverse frozen LLMs. The results of our framework which outperform other prompt-based baselines are indicated with $\uparrow$, the best-performing methods are highlighted in red. L, C, and A denote LLMs of LLaMa2-7B-chat, ChatGPT, and Alpaca, respectively.
@@
RPP (L) 0.82$\pm$0.03 $\uparrow$ 0.85$\pm$0.03$\uparrow$ 0.87$\pm$0.02$\uparrow$ 0.79$\pm$0.03$\uparrow$ 0.82$\pm$0.03$\uparrow$ 0.85$\pm$0.02$\uparrow$ 0.867$\pm$ 0.01$\uparrow$ 0.89$\pm$0.01$\uparrow$ 0.91$\pm$0.01$\uparrow$
@@ 4.3. Overall Results (RQ1)
To answer RQ1, we evaluate the effectiveness of our framework on LLaMa2-7B-chat by comparing the performance of RPP/RPP+ with diverse baselines. We adopt the evaluating metrics of NDCG@1, NDCG@5, and NDCG@10 to access LLMs’ capability in ranking items, the results are listed in Table 2. With task-wise prompts through manual prompts or enumeration, LLaMa2-7B-chat exhibits inferior recommendation performance compared to traditional recommender models and few-shot methods. However, with the instance-wise prompt personalization implemented in the RPP/RPP+, the frozen pre-trained LLaMa2-7B-chat surpasses these baseline models across a variety of datasets. For example, the best-performing baseline in traditional recommender models, SASRec, surpasses the best-performing prompt-based method (enumeration) by $0.64$, $0.48$, and $0.43$ in NDCG@1, NDCG@5, and NDCG@10 metrics, respectively, on the ML-1M dataset. This validates the weakness of task-wise prompts for LLMs. In contrast, RPP demonstrates superior performance compared to the leading baseline in traditional recommender models (SASRec), achieving improvements of $0.14$, $0.01$, and $0.02$ in terms of the NDCG@1, NDCG@5, and NDCG@10 metrics on the ML-1M. Moreover, RPP outperforms the best prompt-based methods (enumeration) with improvements of $0.78$, $0.49$, and $0.45$ regarding the NDCG@1, NDCG@5, and NDCG@10 metrics on the ML-1M. The experimental findings for RPP+ exhibit the same outcomes. This proves the benefit of instance-wise prompts of RPP/RPP+. We conduct multiple tests and the minimal standard deviations demonstrate that the performance improvements resulting from RPP/RPP+ are statistically significant. This illustrates that due to the ability to personalize prompts from instance-wise, RPP/RPP+ enables LLMs to comprehend users’ personalized intents better and provide more accurate recommendations.

### v2

@@ 3.4. Actor-Critic Based Architecture of the Multi-agent and Reward Function
To personalize each pattern individually with the corresponding agent, the Multi-Agent Reinforcement Learning (MARL) is structured with four A2C agents (Peng et al., 2018) under CTDE (Lowe et al., 2017; Sunehag et al., 2018), where each agent has independent parameters but has access to the global state, as shown in Figure 2. The Actor and Critic of each agent are two-layer fully connected neural networks parameterized by $\theta_{g}^{(k)}$ and $\theta_{h}^{(k)}$, denoted as $g^{(k)}(\cdot)$ and $h^{(k)}(\cdot)$, respectively. Specifically, each Actor $g^{(k)}(\cdot)$ makes policy to obtain action $a_{t}^{(k)}$ and its corresponding probability $prob_{t}^{(k)}$ based on the state $\mathbf{s}_{t}$, while each Critic $h^{(k)}(\cdot)$ evaluates the value $v_{t}^{(k)}$ of state $\mathbf{s}_{t}$. Formally, we have:
@@
Since the aim is to boost the recommendation ability of LLMs by personalizing instance-wise prompts, we leverage Normalized Discounted Cumulative Gain (NDCG) directly as the reward, which is the performance metric for ranking tasks. Here we adopt $NDCG@M$ as the reward evaluated according to the answer of LLMs, where $M$ is the number of candidate items and we set it as 10 in our experiments. Formally, given the recommendation output $o_{t}$ at step $t$, we can define the reward $r_{t}$ as below:
@@
(8) $L_{c}^{(k)}$ $=\frac{1}{N}\sum_{t}({\hat{R}_{t-1}}-v_{t}^{(k)}),$
(9) $L_{a}^{(k)}$ $=\frac{1}{N}\sum_{t}\mathrm{log}(prob_{t}^{(k)}({\hat{R}_{t-1}}-v_{t}^{(k)})),$
@@ 3.5. Algorithm
The algorithm for our framework is presented in Algorithm 1.
@@
Algorithm 1 Training Reinforced Prompt Personalization for each epoch.
@@ 4. Experiments
We evaluate the effectiveness of RPP/RPP+ on LLaMa2-7B-chat, comparing it against various baselines including traditional recommender models, few-shot methods, and prompt-based methods. Besides, we extend the evaluation on diverse frozen pre-trained LLMs to measure the RPP/RPP+’s generalization capability. To gain insights into the impact of different patterns in prompts, we conduct an ablation study on personalizing each pattern. Furthermore, we investigate the sensitivity of RPP/RPP+ to hyper-parameters. The case study is employed to provide a tangible illustration of the significance of personalizing prompts in enhancing LLMs’ recommendation ability.
@@ 4.1.1. Datasets
4.1.2. Baselines
We employ several baseline models for comparison, categorizing them into three types: (1) traditional recommender models, including Pop, BPRMF (Rendle et al., 2009), LightGCN (He et al., 2020), and SASRec (Kang and McAuley, 2018); (2) few-shot models, such as BM25 (Robertson and Zaragoza, 2009), VQ-Rec (Hou et al., 2023) and UniSRec (Hou et al., 2022); (3) prompt optimization methods, including manual prompts (Petroni et al., 2019; Schick and Schütze, 2021), enumeration that is a kind of heuristic prompts (Jiang et al., 2020; Gao et al., 2021; Prasad et al., 2023), and GRIPS (Prasad et al., 2023).
Traditional recommender models: Pop relies on item popularity for recommendations but overlooks personalization. BPRMF (Rendle et al., 2009) combines Bayesian methods and matrix factorization techniques to decompose the user-item interaction matrix into a latent factor matrices. LightGCN (He et al., 2020) is a simplified and efficient framework designed for recommendation tasks using graph convolutional networks (GCNs). SASRec (Kang and McAuley, 2018) leverages a self-attention encoder to comprehensively analyze the long-term dependencies present in user behaviors. And it predicts subsequent interactions through a feed-forward network. They are evaluated on the candidate sets after training on the training dataset with all interactions following the setting of LLMRank (Hou et al., 2024)
Prompt-based methods: Manual prompts (Petroni et al., 2019; Schick and Schütze, 2021) are prompts designed by humans to guide LLMs’ behavior, relying on human expertise and domain knowledge. Here we meticulously craft a task-wise prompt for the ranking tasks, encompassing the key patterns within a prompt. Enumeration entails automatically searching within several possible prompts and selecting the best one. Here we design diverse prompts with variations for the ranking tasks, adopting enumeration followed by selection. GRIPS (Prasad et al., 2023) performs edit operations on prompts automatically for LLMs with gradient-free search. Here we adapt it as an instance-wise prompting baseline for ranking tasks.
4.1.3. Frozen Pre-trained LLMs
To evaluate the generalization ability of RPP/RPP+, we extend the experiments on diverse frozen pre-trained LLMs, including LLaMa2, ChatGPT, and Alpaca. They serve as the representative LLMs of three types respectively: the open-source LLM, which can be inferred using pre-trained parameters; the black-box LLM, which solely offers an accessible API; and the fine-tuned LLM, tailored specifically for the recommendation task. LLaMa2 is an open-source LLM released by Meta whose performance has been widely recognized. Here we select the version of LLaMa2-7B-chat as the frozen LLM. ChatGPT is a black-box LLM, which is excellent across a wide array of tasks. We conduct experiments by utilizing the API of GPT-3.5-turbo from OpenAI and set the temperature to 0.2. Alpaca is a lightweight instruction-tuning strategy based on LLaMa. We select LLama2- 7B as the backbone and fine-tune Alpaca with Lora (Hu et al., 2022), which adds trainable rank decomposition matrices to each layer of the transformer architecture in the frozen LLM (Bao et al., 2023), enabling easier and more efficient fine-tuning. We transform the recommendation data for ranking tasks into instruction-tuning data, ensuring that LLMs grasp the specific nuances of the ranking task. The format of the data for instruction tuning can be articulated as follows:
[…]
4.2. Implementation Details
We configure the initial length of interaction history $l_{0}$ as 1 and the number of candidate items to be $M=10$. In our experiments, we set 3, 9, and 5 sentence choices for “role-playing”, “reasoning guidance”, and “output format” patterns in the action spaces respectively. Candidate items are randomly selected from the training dataset and include a ground-truth item sourced from the test dataset. Therefore, for ranking tasks, the HitRatio@10 metric is always 1, rendering this metric meaningless. We set the $\gamma$ in the reward as 0.95. To reduce the randomness of LLMs, we set the temperature as 0.2, since a lower temperature favors more conservative and deterministic outputs. Besides, we report the average performance with corresponding standard deviations of prompt-based methods and RPP/RPP+, based on at least five repeated runs. Traditional recommender models and few-shot-based methods are executed on an NVIDIA RTX 3090Ti, while prompt-based methods and our framework, based on LLMs, run on an NVIDIA A100. During the training phase, we randomly select 200 users as training examples to optimize the multi-agent strategy employed in RPP/RPP+. We terminate the iterative training process if the highest NDCG@10 does not update over 7 iterations or if the total iterations exceed 15. Subsequently, during the testing phase, the multi-agent, now equipped with an optimized policy, undergoes evaluation on a separate set of 100/200 test examples. Our experiments show that on the training set, after three rounds of prompt optimization, the optimized prompts can achieve superior performance on ChatGPT for most instances, while it requires two, seven, and eight iterations on the ML-1M, Games, and Lastfim datasets for LLaMa2-7B-chat, respectively, and approximately eight iterations for Alpaca. Consequently, during the inference process, we terminate the iterations at the corresponding rounds for different datasets and LLMs. We only require users’ collaborative embeddings to initialize the state, along with their interaction history and candidate items as input. To address invalid responses from LLMs, we carefully design prompts to enforce LLMs’ output format, reducing these occurrences. Additionally, we trim and pad the LLMs’ recommendations as 10 while processing invalid responses.
@@ 4.2. Implementation Details
Table 2. Performance comparison of our framework on LLaMa2-7B-chat with baselines across diverse datasets. The results of the best-performing baseline are denoted with an underline. The results of our method which outperform other prompt-based methods are indicated with $\uparrow$, while those that surpass the best-performing baseline are highlighted in red. The performance improvement of RPP/RPP+ compared with the best-performing task-wise prompting method (i.e., Enumeration) is listed in the “improvement”.
@@
Improvement 0.78 0.49 0.45 0.79 0.51 0.49 0.88 0.54 0.48
Relative improvement (%) 1950 136 107 1975 150 132 1760 135 109
@@
Table 3. Additional metrics comparison of our framework on LLaMa2-7B-chat with baselines across diverse datasets, where “M@” means MRR metrics and “H@” means HitRatio metrics. The results of the best-performing baseline are denoted with an underline. The results of our method which outperform other prompt-based methods are indicated with $\uparrow$
methods ML-1M Games Lastfm
M@5 M@10 H@5 M@5 M@10 H@5 M@5 M@10 H@5
SASRec 0.67 0.78 0.80 0.67 0.76 0.87 0.71 0.75 0.81
Manual prompts (task-wise) 0.24$\pm$0.00 0.25$\pm$0.01 0.69$\pm$0.02 0.29$\pm$0.01 0.39$\pm$0.02 0.70$\pm$0.02 0.27$\pm$0.02 0.29$\pm$0.02 0.76$\pm$0.01
Enumeration (task-wise) 0.28$\pm$0.02 0.29$\pm$0.03 0.70$\pm$0.01 0.38$\pm$0.01 0.45$\pm$0.01 0.72$\pm$0.00 0.31$\pm$0.01 0.32$\pm$0.00 0.84$\pm$0.01
GRIPS (instance-wise) 0.52$\pm$0.03 0.54$\pm$0.02 0.74$\pm$0.02 0.51$\pm$0.03 0.52$\pm$0.04 0.73$\pm$0.02 0.50$\pm$0.01 0.51$\pm$0.03 0.76$\pm$0.02
RPP (instance-wise) 0.80$\pm$0.01 $\uparrow$ 0.82$\pm$0.03$\uparrow$ 0.86$\pm$0.03$\uparrow$ 0.80$\pm$0.02$\uparrow$ 0.83$\pm$0.02$\uparrow$ 0.84$\pm$0.03$\uparrow$ 0.85$\pm$ 0.01$\uparrow$ 0.87$\pm$0.02$\uparrow$ 0.87$\pm$0.02$\uparrow$
RPP+ (instance-wise) 0.82$\pm$0.02 $\uparrow$ 0.83$\pm$0.01$\uparrow$ 0.87$\pm$0.00$\uparrow$ 0.85$\pm$0.01$\uparrow$ 0.86$\pm$0.00$\uparrow$ 0.87$\pm$0.01$\uparrow$ 0.88$\pm$ 0.01$\uparrow$ 0.89$\pm$0.00$\uparrow$ 0.89$\pm$0.01$\uparrow$
Table 4. Performance of RPP and RPP+ on diverse frozen LLMs. The results of our framework which outperform other prompt-based baselines are indicated with $\uparrow$, and the best-performing methods are highlighted in red. L, C, and A denote LLMs of LLaMa2-7B-chat, ChatGPT, and Alpaca, respectively.
@@
GRIPS (L) 0.43$\pm$0.01 0.70$\pm$0.03 0.73$\pm$0.05 0.62$\pm$0.04 0.67$\pm$0.02 0.72$\pm$0.02 0.19$\pm$0.03 0.64$\pm$0.01 0.66$\pm$0.04
RPP (L) 0.82$\pm$0.03 $\uparrow$ 0.85$\pm$0.03$\uparrow$ 0.87$\pm$0.02$\uparrow$ 0.79$\pm$0.03$\uparrow$ 0.82$\pm$0.03$\uparrow$ 0.85$\pm$0.02$\uparrow$ 0.87$\pm$ 0.01$\uparrow$ 0.89$\pm$0.01$\uparrow$ 0.91$\pm$0.01$\uparrow$
@@
Table 7. Additional experiments for the pattern of “user profile”.
methods ML-1M
N@1 N@5 N@10
Manual prompts (task-wise) 0.05$\pm$0.01 0.41$\pm$0.02 0.46$\pm$0.01
Enumeration (task-wise) 0.16$\pm$0.00 0.48$\pm$0.02 0.51$\pm$0.02
GRIPS (instance-wise) 0.63$\pm$0.01 0.72$\pm$0.02 0.77$\pm$0.00
user profile (instance-wise) 0.78$\pm$0.01 $\uparrow$ 0.79$\pm$0.03$\uparrow$ 0.82$\pm$0.03$\uparrow$
RPP (instance-wise) 0.83$\pm$0.02 $\uparrow$ 0.83$\pm$0.01$\uparrow$ 0.88$\pm$0.01$\uparrow$
RPP+ (instance-wise) $0$uparrow 0.84$\pm$0.01$\uparrow$ 0.89$\pm$0.00$\uparrow$
Table 8. Ablation study on the impact of GRU.
methods ML-1M Games Lastfm
N@1 N@5 N@10 N@1 N@5 N@10 N@1 N@5 N@10
GRIPS 0.43$\pm$0.01 0.70$\pm$0.03 0.73$\pm$0.05 0.62$\pm$0.04 0.67$\pm$0.02 0.72$\pm$0.02 0.19$\pm$0.03 0.64$\pm$0.01 0.66$\pm$0.04
RPP-pool 0.80$\pm$0.00 0.84$\pm$0.01 0.86$\pm$0.02 0.76$\pm$0.03 0.79$\pm$0.00 0.81$\pm$0.02 0.90$\pm$ 0.01 0.92$\pm$0.01 0.93$\pm$0.00
RPP-Bert 0.82$\pm$0.01 0.85$\pm$0.00 0.87$\pm$0.01 0.83$\pm$0.02 0.85$\pm$0.01 0.86$\pm$0.01 0.93$\pm$ 0.02 0.94$\pm$0.01 0.95$\pm$0.02
@@ 4.3. Overall Results (RQ1)
To answer RQ1, we evaluate the effectiveness of our framework on LLaMa2-7B-chat by comparing the performance of RPP/RPP+ with diverse baselines. We adopt the evaluating metrics of NDCG@1, NDCG@5, and NDCG@10 to access LLMs’ capability in ranking items, the results are listed in Table 2. Our observations are as follows: 1) Task-wise prompting methods (i.e., manual prompts and enumeration) exhibit inferior recommendation performance compared to traditional recommender models, few-shot methods, and instance-wise prompting methods (i.e., GRIPS and RPP/RPP+). For example, the best-performing baseline in traditional recommender models, SASRec, surpasses the best-performing task-wise prompting method (enumeration) by $0.64$, $0.48$, and $0.43$ in NDCG@1, NDCG@5, and NDCG@10 metrics, respectively, on the ML-1M dataset. This validates the weakness of task-wise prompts for LLMs. 2) Instance-wise prompting implemented by RPP/RPP+ makes the frozen pre-trained LLaMa2-7B-chat surpass traditional recommender models across a variety of datasets. For instance, RPP demonstrates superior performance compared to the leading baseline in traditional recommender models (SASRec), achieving improvements of $0.14$, $0.01$, and $0.02$ in terms of the NDCG@1, NDCG@5, and NDCG@10 metrics on the ML-1M. 3) The framework of RPP/RPP+ outperforms other instance-wise promoting methods in the ranking tasks, such as GRIPS. This proves the high quality of prompts generated by RPP/RPP+. We conduct multiple tests and the minimal standard deviations demonstrate that the performance improvements resulting from RPP/RPP+ are statistically significant. Additionally, we adopt the MRR@5, MRR@10, and HitRatio@5 metrics to evaluate the performance of RPP/RPP+ in Table 3, as these are widely used benchmarks. The results provide valuable insights that highlight the superiority of our method, which achieves state-of-the-art performance compared to the baselines. These overall results illustrate that due to the ability to personalize prompts from instance-wise, RPP/RPP+ enables LLMs to comprehend users’ personalized intents better and provide more accurate recommendations.